OpenStudy (anonymous):

Find the exact value of sin^2 (pi/8). Half angle identity may be used.

OpenStudy (xapproachesinfinity):

cos2x=1-2sin^2x sin^2(x)=(1+cos2x)/2 x=pi/8 sin^2(pi/8)=(1+cos(pi/4)/2=(1+root2/2)/2